Determine the mass of CO2 produced by burning enough carbon (in the form of charcoal) to produce 4.50×10^2 kJ of heat.

C(s)+O2(g)→CO2(g)

ΔrH∘= −393.5kJ mol−1
Answer #1

C(s)+O2(g)→CO2(g) ΔrH∘= −393.5kJ mol−1

One mole of CO2 produces from one mole of C and gives -393.5 kJ heat.

Now x moles of CO2 gives 450 kJ of heat.

x moles of CO2 = 450/393.5 = 1.143 moles of CO2

Mass of 1.143 moles of CO2 = 1.143 × 44

= 50.31 g of CO2 produces